OpenStudy (anonymous):

Find the missing coordinate. (-3,4) is a solution of y=4-3x

OpenStudy (anonymous):

I don't understand your question of finding the missing coordinate. However, for a point like (-3,4) to be a solution of an equation it must make a true statement. For example: if we have the equation y=3x, (1,3) is a solution to the equation because when x=1, y=3 it makes a true statement for: y=3x 3=3(1) 3=3 this is true. But, (-3,4) is not a solution of y=4-3x because when x=-3 and y=4 we have y=4-3x 4=4-3(-3) 4=4+9 4=13 is a FALSE STATEMENT hope this helps

OpenStudy (anonymous):

thank you so much it help a lot.